# Fluctuations and the limit of predictability in protein evolution

### v1

@@ Abstract
Protein evolution involves mutations occurring across a wide range of time scales [Di Bari et al., PNAS 121, e2406807121 (2024)]. In analogy with other disordered systems, this dynamical heterogeneity suggests strong correlations between mutations happening at distinct sites and times. To quantify these correlations, we examine the role of various fluctuation sources in protein evolution, simulated using a data-driven epistatic landscape. By applying spatio-temporal correlation functions inspired by statistical physics, we disentangle fluctuations originating from the ancestral protein sequence from those driven by stochastic mutations along independent evolutionary paths. Our analysis shows that, in diverse protein families, fluctuations from the ancestral sequence predominate at shorter time scales. This allows us to identify a time scale over which ancestral sequence information persists, enabling its reconstruction. We link this persistence to the strength of epistatic interactions: ancestral sequences with stronger epistatic signatures impact evolutionary trajectories over extended periods. At longer time scales, however, ancestral influence fades as epistatically constrained sites evolve collectively. To confirm this idea, we apply a standard ancestral sequence reconstruction algorithm and verify that the time-dependent recovery error is influenced by the properties of the ancestor itself.
@@ I Introduction
Understanding and characterizing the impact of epistasis in evolution requires careful experiments and modeling. In particular, recent developments increased substantially the sequence divergence that can be reached by laboratory evolution experiments Fantini et al. (2020); Stiffler et al. (2020); Erdoğan et al. (2023); Park et al. (2022); Rix et al. (2023), such that many data are becoming available and many more are expected to come soon. Yet, the sequence diversity of natural evolution still remains out of reach of such experiments, thus leaving an unexplored gap in evolutionary time scales. In order to fill this gap, one can simulate the evolution of protein sequences in silico de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2024); Biswas et al. (2024); Bari et al. (2024), relying on the data-driven approach that goes under the name of Direct Coupling Analysis (DCA) Morcos et al. (2011); Cocco et al. (2018), in which a fitness landscape (analogous to an energy function in the statistical physics vocabulary) is inferred starting from a Multiple Sequence Alignment (MSA) of natural homologs constituting a given protein family Ferguson et al. (2013); Figliuzzi et al. (2016); Levy et al. (2017); Couce et al. (2017); Vigué and Tenaillon (2023). The energy function that results from this inference procedure can then be used to assign a probability to each sequence. The resulting landscape is explored by means of a biologically motivated Monte Carlo Markov-Chain (MCMC) algorithm, providing in silico evolutionary trajectories that quantitatively mimic experimental results de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2022); Biswas et al. (2024); Bari et al. (2024). Within this framework, it has been recently shown Bari et al. (2024) that different sites evolve with widely different time scales, which also depend on the background sequence, due to epistatic interactions. By analogy with similar phenomena observed in disordered physical systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018), this large heterogeneity in time scales suggests the presence of strong and highly collective correlations between sites that slow down the evolution of the sequence. This is the main observation that motivated this work.
@@
In this paper, we measure such correlations to explore how the heterogeneity of evolutionary time scales relates to the epistatic interactions within the ancestral sequence. We follow the dynamics of the Hamming distance, i.e., the number of accepted mutations with respect to the ancestral sequence. We characterize how this quantity fluctuates (i) between different evolutionary trajectories originating from the same ancestor, and (ii) between different ancestors. We observe a strong dependence of the dynamics on the ancestral sequence for short enough times, over which the first source of fluctuations is found to be subdominant with respect to the second. On the other hand, at long times the stochasticity of evolutionary trajectories takes over, and any memory of the ancestor is lost. This has important consequences, as it allows us to properly quantify the time scale over which it should be possible to reconstruct the ancestral sequence of a certain set of evolutionary trajectories, and how this time scale depends on the epistatic interactions in the ancestor itself. In fact, we find that the amount of epistatically constrained sites in the ancestral sequence determines this time scale: more epistatic sequences leave their trace on evolutionary dynamics for longer times. We then measure the correlations between the evolution of all pairs of sites at the time scale at which such correlations are stronger, finding a different pattern for each ancestral sequence, that is then reflected in the evolutionary dynamics. Finally, we show that the magnitude of epistatic interactions, which controls the stochasticity of the evolutionary trajectories, is also related to the linear response of the evolutionary dynamics to a change in selective pressure (controlled by the temperature in our Monte Carlo simulations). Hence, we show and quantify how more epistatically constrained ancestors lead to a more complex evolutionary dynamics over longer time scales, which is also more sensitive to perturbations of the environment.
@@ II.3 Dynamical fluctuations
Building on the previous work described above, and inspired by works on disordered systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018), we devised a protocol (sketched in Fig. 1) to quantify how the amount of epistasis in the ancestral sequence impacts the subsequent evolution. We consider a series of ancestral sequences, taken from the natural MSA, let us call them $A_{0}$, $B_{0}$, $C_{0}$, etc. Each ancestor has a different level of epistatic constraints, as measured by the average CDE introduced in Eq. (4), as illustrated in Fig. 1a.
@@
To make this observation quantitative, we characterize the statistical properties of the resulting MSAs, and how they depend on the ancestor, by introducing two distinct averages and corresponding fluctuations, inspired from the disordered systems literature Franz et al. (2011, 2013); Seoane and Zamponi (2018) and illustrated in Fig. 1b. Recall that the number of accepted mutations $H^{A_{0}}(t)$ at evolutionary time $t$ for fixed ancestor $A_{0}$ is a random variabile, whose realizations depend on the stochasticity of the evolutionary trajectory.
@@ III.6 Limits of ancestral sequence reconstruction
In an evolutionary process, predictions can be made in two directions: either starting from the ancestral sequence and predicting its future evolution over a given time; or starting from a set of evolved sequences and attempting to infer the ancestral sequence from which the evolution began. An attempt at predicting future evolution has been made in Ref. Bisardi et al. (2022), where however the prediction was limited to small evolutionary time scales. Because we have repeatedly hinted that our analysis can provide insight into the time scale over which reconstructing the ancestral sequence of an evolutionary process is possible, before concluding, we consider here the ancestral sequence reconstruction problem more explicitly.
@@ IV Conclusions
In this study, we explore the role of disorder and stochasticity in protein evolution, focusing on the interplay between epistasis, site correlations, and sequence-dependent fluctuations. We simulated evolution in silico using DCA to model the fitness landscape, and a MCMC dynamics to mimic the process of mutation and selection, a methodology that has been validated in previous work de la Paz et al. (2020); Bisardi et al. (2022); Bari et al. (2024). By employing tools from statistical physics Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018), we quantified how the ancestral amino acid sequence significantly influences early evolutionary dynamics, as measured by the dynamical susceptibility. Our main results are the following.

### v2

@@ Abstract
Protein evolution involves mutations occurring across a wide range of time scales. In analogy with other disordered systems, this dynamical heterogeneity suggests strong correlations between mutations happening at distinct sites and times. To quantify these correlations, we examine the role of various fluctuation sources in protein evolution, simulated using a data-driven epistatic landscape. By applying spatio-temporal correlation functions inspired by statistical physics, we disentangle fluctuations originating from the ancestral protein sequence from those driven by stochastic mutations along independent evolutionary paths. Our analysis shows that, in diverse protein families, fluctuations from the ancestral sequence predominate at shorter time scales. This allows us to identify a time scale over which ancestral sequence information persists, enabling its reconstruction. We link this persistence to the strength of epistatic interactions: ancestral sequences with stronger epistatic signatures impact evolutionary trajectories over extended periods. At longer time scales, however, ancestral influence fades as epistatically constrained sites evolve collectively. To confirm this idea, we apply a standard ancestral sequence reconstruction algorithm and verify that the time-dependent recovery error is influenced by the properties of the ancestor itself.
@@ I Introduction
Understanding and characterizing the impact of epistasis in evolution requires careful experiments and modeling. In particular, recent developments increased substantially the sequence divergence that can be reached by laboratory evolution experiments Fantini et al. (2020); Stiffler et al. (2020); Erdoğan et al. (2023); Park et al. (2022); Rix et al. (2023), such that many data are becoming available and many more are expected to come soon. Yet, the sequence diversity of natural evolution still remains out of reach of such experiments, thus leaving an unexplored gap in evolutionary time scales. In order to fill this gap, one can simulate the evolution of protein sequences in silico de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2024); Biswas et al. (2024); Bari et al. (2024), relying on the data-driven approach that goes under the name of Direct Coupling Analysis (DCA) Morcos et al. (2011); Cocco et al. (2018), in which a fitness landscape (analogous to an energy function in the statistical physics vocabulary) is inferred starting from a Multiple Sequence Alignment (MSA) of natural homologs constituting a given protein family Ferguson et al. (2013); Figliuzzi et al. (2016); Levy et al. (2017); Couce et al. (2017); Vigué and Tenaillon (2023). The energy function that results from this inference procedure can then be used to assign a probability to each sequence. The resulting landscape is explored by means of a biologically motivated Monte Carlo Markov-Chain (MCMC) algorithm, providing in silico evolutionary trajectories that quantitatively mimic experimental results de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2022); Biswas et al. (2024); Bari et al. (2024).
Within this framework, it has been recently shown Bari et al. (2024) that different sites evolve with widely different time scales, which also depend on the background sequence, due to epistatic interactions. This heterogeneity of local time scales is strongly reminiscent of similar phenomena observed in disordered physical systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Berthier and Jack (2007); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018). There, it has been shown that a large heterogeneity in time scales implies the presence of strong and highly collective space-time correlations between sites that slow down the dynamical evolution. Furthermore, it has been shown that the structural disorder in the initial configuration partially encodes future correlations, and can be used (possibly by machine learning tools) to predict the future dynamics Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2024), see Jung et al. (2023) for a recent review. These are the main observations that motivated this work.
@@
In this paper, we measure space-time correlations to explore how the heterogeneity of evolutionary time scales relates to the epistatic interactions within the ancestral sequence. We follow the dynamics of the Hamming distance, i.e., the number of accepted mutations with respect to the ancestral sequence. We characterize how this quantity fluctuates (i) between different evolutionary trajectories originating from the same ancestor, and (ii) between different ancestors. We observe a strong dependence of the dynamics on the ancestral sequence for short enough times, over which the first source of fluctuations is found to be subdominant with respect to the second. On the other hand, at long times the stochasticity of evolutionary trajectories takes over, and any memory of the ancestor is lost. This has important consequences, as it allows us to properly quantify the time scale over which it should be possible to reconstruct the ancestral sequence of a certain set of evolutionary trajectories, and how this time scale depends on the epistatic interactions in the ancestor itself. In fact, we find that the amount of epistatically constrained sites in the ancestral sequence determines this time scale: more epistatic sequences leave their trace on evolutionary dynamics for longer times. We then measure the correlations between the evolution of all pairs of sites at the time scale at which such correlations are stronger, finding a different pattern for each ancestral sequence, that is then reflected in the evolutionary dynamics. Finally, we show that the magnitude of epistatic interactions, which controls the stochasticity of the evolutionary trajectories, is also related to the linear response of the evolutionary dynamics to a change in selective pressure (controlled by the temperature in our Monte Carlo simulations). Hence, we show and quantify how more epistatically constrained ancestors lead to a more complex evolutionary dynamics over longer time scales, which is also more sensitive to perturbations of the environment.
@@ II.3 Dynamical fluctuations
Building on the previous work described above, and inspired by works on disordered systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018); Berthier and Jack (2007), we devised a protocol (sketched in Fig. 1) to quantify how the amount of epistasis in the ancestral sequence impacts the subsequent evolution. We consider a series of ancestral sequences, taken from the natural MSA, let us call them $A_{0}$, $B_{0}$, $C_{0}$, etc. Each ancestor has a different level of epistatic constraints, as measured by the average CDE introduced in Eq. (4), as illustrated in Fig. 1a.
@@
To make this observation quantitative, we characterize the statistical properties of the resulting MSAs, and how they depend on the ancestor, by introducing two distinct averages and corresponding fluctuations, inspired from the disordered systems literature Franz et al. (2011, 2013); Seoane and Zamponi (2018); Berthier and Jack (2007) and illustrated in Fig. 1b. Recall that the number of accepted mutations $H^{A_{0}}(t)$ at evolutionary time $t$ for fixed ancestor $A_{0}$ is a random variabile, whose realizations depend on the stochasticity of the evolutionary trajectory.
@@ III.6 Limits of ancestral sequence reconstruction
In an evolutionary process, predictions can be made in two directions: either starting from the ancestral sequence and predicting its future evolution over a given time; or starting from a set of evolved sequences and attempting to infer the ancestral sequence from which the evolution began. An attempt at predicting future evolution has been made in Ref. Bisardi et al. (2022) (see Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2023, 2024) for related studies in disordered systems), where however the prediction was limited to small evolutionary time scales. Because we have repeatedly hinted that our analysis can provide insight into the time scale over which reconstructing the ancestral sequence of an evolutionary process is possible, before concluding, we consider here the ancestral sequence reconstruction problem more explicitly.
@@ IV Conclusions
In this study, we explore the role of disorder and stochasticity in protein evolution, focusing on the interplay between epistasis, site correlations, and sequence-dependent fluctuations. We simulated evolution in silico using DCA to model the fitness landscape, and a MCMC dynamics to mimic the process of mutation and selection, a methodology that has been validated in previous work de la Paz et al. (2020); Bisardi et al. (2022); Bari et al. (2024). By employing tools from statistical physics Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018); Berthier and Jack (2007), we quantified how the ancestral amino acid sequence significantly influences early evolutionary dynamics, as measured by the dynamical susceptibility. Our main results are the following.